qweqeeqw:

Two man who are 475 m apart, find the angles of elevation of a kite between them are 28 degree and 42 degree, respectively. what is the altitude of the kite, which is the same plane as them (math problem)

jhonyy9:

try draw it

qweqeeqw:

can you also give a solution thankyou so much

jhonyy9:



qweqeeqw:

thankyou so much can you also give a solution? that would be really help

jhonyy9:

the altitude of the kite let be h on the laft part let be the man 1 and on the right part man 2 the distance between 1 man and the kite let be x and the distance between 2 man let be y using these you can write tan 28° = h/(475-y) tan 42° = h/(475-x) any idea now ?

jhonyy9:

or use @florisalreadytaken's idea with the law of sin

qweqeeqw:

thankyou appreciated

Florisalreadytaken:

so, law of sines tells us that: \[\frac{\sin \alpha}{a}=\frac{\sin \gamma}{c} \] look at the image i attached below, for the namng of the things.

1 attachment
Florisalreadytaken:

\[\frac{\sin 42^o }{a}=\frac{\sin 110^o }{475} \Rightarrow a \sin \left(110^o \right)=475 \sin \left(42^o \right) \Rightarrow a=\frac{475 \sin \left(42^o \right) }{\sin \left(110^o\right)} \] \[ a=?? \]

qweqeeqw:

181?

Florisalreadytaken:

no. \[ a \approx 338.24 \] the distance from the 1st guy to the kite is 338.24 m.

qweqeeqw:

thankyou

Florisalreadytaken:

by using SOH CAH TOA, we get: \[ \sin(28^o) = \frac{opposite}{hypotenuse} \] \[ \sin(28^o) = \frac{h}{a} \Rightarrow \sin(28^o) = \frac{h}{338.24} \] \[ h=338.24 \sin(28^o) \] \[ \Large h=?? \]

qweqeeqw:

159

Florisalreadytaken:

no... \[ \Large h=158.63 \] and that's tour answer

qweqeeqw:

so thats the final ans?

Florisalreadytaken:

positive

qweqeeqw:

okay thankyou so much